### v2

# GLEAM: Learning to Match and Explain in Cross-View Geo-Localization

Abstract
Cross-View Geo-Localization (CVGL) focuses on identifying correspondences between images captured from distinct perspectives of the same geographical location. However, existing CVGL approaches are typically restricted to a single view or modality, and their direct visual matching strategy lacks interpretability: they only determine whether two images correspond, without explaining the rationale behind the match. In this paper, we present GLEAM-C, a foundational CVGL model that unifies multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—by aligning them exclusively with satellite imagery. Our framework enhances training efficiency through optimized implementation while achieving accuracy comparable to prior modality-specific CVGL models through a two-phase training strategy. Moreover, to address the lack of interpretability in traditional CVGL methods, we leverage the reasoning capabilities of multimodal large language models (MLLMs) to propose a new task, GLEAM-X, which combines cross-view correspondence prediction with explainable reasoning. To support this task, we construct a bilingual benchmark using GPT-4o and Doubao-1.5-Thinking-Vision-Pro to generate training and testing data. The test set is further refined through detailed human revision, enabling systematic evaluation of explainable cross-view reasoning and advancing transparency and scalability in geo-localization. Together, GLEAM-C and GLEAM-X form a comprehensive CVGL pipeline that integrates multi-modal, multi-view alignment with interpretable correspondence analysis, unifying accurate cross-view matching with explainable reasoning and advancing Geo-Localization by enabling models to better Explain And Match. Code and datasets used in this work will be made publicly accessible at https://github.com/Lucky-Lance/GLEAM.
1 Introduction
Cross-View Geo-Localization (CVGL) seeks to determine the geographic position of a query image by establishing correspondences with a geo-referenced database captured from alternative viewpoints, such as satellite observations. Unlike GNSS, which may suffer from signal blockage or noise in urban canyons and complex environments, CVGL offers a vision-based localization paradigm that is both complementary and robust. This capability has broad relevance to autonomous driving (Cui & Ge, 2003; Chen et al., 2023), robotic navigation (Nowak et al., 2024; Semborski & Idzkowski, 2024), unmanned aerial vehicles (UAVs) navigation (Xu et al., 2024; Suzuki et al., 2016; Wang et al., 2024), and augmented reality devices (Kamalam et al., 2022; Sathyanarayana et al., 2020), where reliable positioning forms the foundation for effective operation.
Despite its practical importance, CVGL remains a technically demanding problem. The task is complicated by the substantial appearance gap across heterogeneous viewpoints and modalities (Regmi & Shah, 2019; Ge et al., 2024). Consequently, prior studies have typically focused on a single viewpoint and modality, such as UAV imagery (Zheng et al., 2020), street maps (Zhao et al., 2021), panoramic views (Zhu et al., 2021), or ground photographs (Wu et al., 2024b). While these modality-specific approaches have achieved promising results, the problem of establishing robust cross-modal alignment within a unified framework remains unresolved. A unified CVGL framework offers several advantages. First, it enables multi-platform deployment through a single inference interface, reducing system complexity, memory footprint, and initialization latency on resource-constrained platforms such as drones and vehicles. Second, it improves robustness and generalization by learning a geographically grounded representation that captures diverse structural regularities. By bridging heterogeneous views and modalities, such a system enables robust, generalizable, and efficient geo-positioning across diverse real-world scenarios.
Figure 1: Overview of GLEAM-Core and GLEAM-eXplain. (A) GLEAM-C: a foundational CVGL model trained to align multiple views and modalities—including UAV imagery, street maps, panoramic images, and ground photos—with satellite imagery. (B) GLEAM-X: a benchmark combining cross-view correspondence prediction with explainable reasoning. We illustrate representative examples of both positive and negative matches between query images and satellite images. The red arrow indicates that GLEAM-C and GLEAM-X can be combined into an integrated pipeline.
Moreover, the dominant paradigm in CVGL has been image retrieval or binary correspondence prediction (Deuser et al., 2023a; Xia et al., 2024b), which offers limited interpretability and constrains system transparency. For safety-critical applications such as navigation or disaster response, a mere correspondence score is insufficient; instead, models must provide interpretable reasoning about why two views are matched or mismatched. Recent advances in multimodal large language models (MLLMs) present an avenue for addressing this limitation. By jointly processing visual and textual information, MLLMs can reason about image correspondences and generate human-interpretable explanations (Achiam et al., 2023; Bai et al., 2025a; Zhu et al., 2025). Integrating these reasoning capabilities into CVGL enables models not only to predict whether two images correspond but also to articulate the underlying rationale, enhancing transparency and accountability in decision-making.
To address these two challenges, in this paper, we propose a comprehensive solution that combines robust multi-modal alignment with interpretable cross-view reasoning. The overview of our work is shown in Fig. 1. First, we design GLEAM-Core (GLEAM-C), a foundational CVGL model that integrates multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—by aligning them exclusively with satellite images. This unified design eliminates the need for modality-specific architectures, thereby simplifying the modeling process and improving scalability across heterogeneous data sources. In GLEAM-C, we reconstruct the conventional Data Parallel (DP) training strategy used in prior work (Deuser et al., 2023b) into a Distributed Data Parallel (DDP) scheme, which yields over 5-fold improvement in training efficiency. We design a two-phase training strategy and evaluate GLEAM-C on both CNN- and ViT-based architectures. The results show accuracy comparable to that of modality-specific models. Second, we design GLEAM-eXplain (GLEAM-X), which combines cross-view correspondence prediction with explainable reasoning. To enhance model robustness, we select 50k query images from the multi-view matching training set, each paired with a positive reference image (match) and a negative reference image (mismatch). After assigning ground-truth labels, GPT-4o and Doubao-1.5-Thinking-Vision-Pro generate explanations in both Chinese and English, clarifying the reasons for image correspondences. For evaluation, we select 504 query images from the matching test set, each paired with a positive and a negative reference image. After GPT-4o generates explanations, we perform multiple rounds of human annotation and refinement to create a high-quality test set. We fine-tune the Qwen2.5-VL-3B-Instruct MLLM on the training set, achieving higher matching accuracy than GPT-4o and Doubao-1.5, with analysis results closely aligned with human annotations. GLEAM-C and GLEAM-X can be further integrated: GLEAM-C performs the core task of cross-view geo-localization by aligning query images with geo-referenced satellite imagery. GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the image correspondences predicted by GLEAM-C, improving both the model’s robustness and transparency. The contributions of our work can be summarized as follows:
1) Problem Analysis: We provide a detailed analysis of the CVGL task, highlight the benefits of a unified CVGL framework, and identify the lack of interpretability in current CVGL solutions.
2) GLEAM-C: We propose GLEAM-C, a foundational CVGL model that unifies multiple views and modalities—including UAV images, street maps, panoramic images, and ground photographs—by aligning them with satellite images. We adopt a two-phase training strategy and leverage DDP to optimize training efficiency, achieving accuracy comparable to modality-specific CVGL models.
3) GLEAM-X: We introduce the GLEAM-X benchmark, combining cross-view correspondence prediction with explainable reasoning. We provide a high-quality bilingual (Chinese-English) training and testing dataset, along with the Qwen2.5-VL-3B-Instruct model fine-tuned on this dataset. This enables the generation of human-interpretable explanations for image correspondences, thereby enhancing the robustness and transparency of the model.
4) Integrated Pipeline: GLEAM-C and GLEAM-X can be combined into an integrated pipeline. GLEAM-C performs the core task of CVGL by aligning query images with satellite imagery, while GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the predicted correspondences, improving both the model’s robustness and transparency.
2 Related Works
2.1 Cross-View Geo-Localization
Cross-View Geo-Localization (CVGL) has garnered substantial attention within the research community. Most existing research, however, focuses on a single view or modality. Early efforts construct paired ground-to-aerial datasets (Lin et al., 2015), which later evolve into widely used benchmarks such as CVUSA (Workman et al., 2015), CVACT (Liu & Li, 2019), VIGOR (Zhu et al., 2021), University-1652 (Zheng et al., 2020), and DenseUAV (Dai et al., 2023), enabling evaluations across panoramas, UAV imagery, and related scenarios. Methodologically, handcrafted descriptors (Bansal et al., 2011; Castaldo et al., 2015) soon give way to deep learning, where pre-trained CNNs and fine-tuning strategies significantly advance cross-view correspondence (Krizhevsky et al., 2012; Workman et al., 2015). Subsequent studies introduce a variety of techniques to alleviate viewpoint discrepancies, including polar and optimal transport transformations (Shi et al., 2020; 2019), region-level and latent alignment (Dai et al., 2021; Xia et al., 2024a), and strategies that enhance scene discrimination, such as hard-negative mining (Deuser et al., 2023b) and cross-dimension interactions (Shen et al., 2023). More recently, unsupervised approaches explore how to exploit unlabeled data for training (Li et al., 2024c; b). Despite these advancements, most studies remain modality-specific, and the development of a unified framework that accommodates diverse modalities still represents an open challenge.
2.2 Multimodal Large Language Model
Multimodal large language models (MLLMs) have emerged in recent years as a powerful tool for handling complex tasks by incorporating visual inputs into traditional language models (OpenAI, 2023; Liu et al., 2024; Chen et al., 2024; Bai et al., 2025a). Prominent commercial models such as OpenAI’s GPT-4o (Hurst et al., 2024) and Google’s Gemini 2.5 (Comanici et al., 2025) exemplify this trend by simultaneously handling text, audio, images, and video, unlocking new capabilities like real-time, human-like voice interaction and long-context video understanding. In addition to closed-source commercial MLLMs, numerous well-known open-source models, such as InternVL3 (Wang et al., 2025), Qwen2.5-VL (Bai et al., 2025b), MiniCPM-V 4.5 (Yao et al., 2024), and LLaVA-OneVision (Li et al., 2024a), have made significant contributions to advancing multimodal understanding. Beyond perception alone, recent efforts increasingly emphasize unifying multimodal understanding and generation, thereby enabling models not only to interpret visual inputs but also to synthesize new content. Representative examples include Janus-Pro (Chen et al., 2025), Show-o (Xie et al., 2024), and BAGEL (Deng et al., 2025), which collectively mark a transition toward more general-purpose multimodal intelligence that integrates both reasoning and creativity.
Figure 2: Method Overview of GLEAM-C and GLEAM-X. (A) GLEAM-C: We apply a contrastive learning architecture to train the CVGL model across UAV, street map, panoramic, and ground photographs. (B) GLEAM-X: This component formulates a multi-image reasoning task. The MLLM receives a query image, a reference image, and a natural language instruction. Through fine-tuning, it delivers both a matching prediction and an interpretable textual explanation.
3 GLEAM-C: A Unified CVGL Core Model
GLEAM-C is designed as a unified CVGL model, leveraging a contrastive learning paradigm to align visual representations across diverse viewpoints and modalities. In this section, we provide the data composition (Sec. 3.1), model structure (Sec. 3.2), and training recipe (Sec. 3.3).
3.1 Data Composition
Dataset Original Pairs GLEAM-C GLEAM-X
Sampling Ratio Training Pairs Sampling Ratio
Training Pairs
(pos+neg, en+zh)
University-1652 (Zheng et al., 2020) 37,854 1.00 37,854 0.36 54,000
VIGOR (Zhu et al., 2021) 52,609 1.00 52,609 0.26 54,000
SetVL-480K (Wu et al., 2024a) 240,544 0.50 120,272 0.06 54,000
MAP 10,208 4.00 40,832 1.00 40,832
Total 341,215 251,567 202,832
Table 1: Statistics of training pairs in GLEAM-C and GLEAM-X. GLEAM-C primarily constructs training pairs from existing CVGL datasets. GLEAM-X balances the data distribution across datasets and then samples positive and negative pairs, providing bilingual explanation annotations.
GLEAM-C is a foundational CVGL model that aligns multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—exclusively with satellite imagery. To facilitate training, we utilize widely adopted datasets from prior work: University-1652 (Zheng et al., 2020) for UAV imagery, VIGOR (Zhu et al., 2021) for panoramic views, and SetVL-480K (Wu et al., 2024a) for ground photographs. In addition, for street maps, we manually collect 12,761 pairs of corresponding street maps and satellite images from Google Maps, with each image resized to $\times512$ pixels. We use 10,208 pairs for training and 2,553 pairs for testing.
The original datasets show highly imbalanced sample distributions across different views and modalities, as shown in Tab. 1, which could introduce biases and hinder the model from learning uniformly. To mitigate this issue, we employ data sampling to balance the number of samples across all views and modalities. This ensures that the model receives sufficient and diverse training signals from each type of data, thereby improving its robustness across heterogeneous inputs. For evaluation, we adhere to the original test datasets and protocols provided by each benchmark. The newly introduced MAP test follows the same evaluation logic as VIGOR (Zhu et al., 2021).
3.2 Model Structure
We leverage a contrastive training strategy for GLEAM-C, where the query and reference images share the same encoder for feature extraction. The InfoNCE loss is applied to learn discriminative features across both view directions. The training framework is shown in Fig. 2A. Our implementation is based on the well-engineered Sample4Geo (Deuser et al., 2023b) codebase. Notably, Sample4Geo introduces GPS-Sampling (GPS) and Dynamic Similarity Sampling (DSS) to select hard negatives during training, thereby improving optimization efficiency. However, certain datasets, such as University-1652 (Zheng et al., 2020), do not provide GPS metadata. Therefore, GLEAM-C relies solely on DSS. In our experiments, we evaluate both convolutional and transformer-based architectures as backbones. Specifically, we adopt ConvNeXt (Liu et al., 2022), comprising 88.6M parameters, and the Perception Encoder (Bolya et al., 2025), comprising 0.32B parameters.
3.3 Training Recipe
We propose a two-phase training scheme to effectively integrate knowledge across multiple modalities and views (Sec. 3.3.1). Given the substantial data requirements of foundational model training and the limited efficiency of the original Sample4Geo codebase, we implement targeted optimizations to significantly enhance training efficiency (Sec. 3.3.2).
3.3.1 Two-Phase Training Strategy
We note that a naive concatenation of multiple datasets for training from the outset poses significant challenges. Due to the backbone model’s limited matching capacity and the varying dataset sizes and difficulty levels, direct multi-dataset training often results in highly imbalanced performance. To mitigate this issue, we first train the model on a single dataset that is relatively large and of moderate difficulty, allowing the backbone to acquire fundamental CVGL capabilities before introducing more complex and diverse data. In our design, we first train the model on the VIGOR (Zhu et al., 2021) dataset for 40 epochs, followed by training on the concatenated datasets for an additional 40 epochs.
3.3.2 Training Efficiency Improvement
Similar to CLIP training, contrastive learning models typically require large batch sizes. The original Sample4Geo adopts PyTorch’s Data Parallel (DP) strategy, where GPU 0 distributes data, aggregates outputs, and performs gradient updates for the entire cluster. This centralized design creates a significant communication bottleneck on GPU 0, thus slowing the training process. Distributed Data Parallel (DDP) alleviates this issue by allowing each GPU to compute gradients locally and synchronizing them via an efficient all-reduce. However, contrastive learning relies on in-batch negatives, and small per-GPU batch sizes limit the number of negatives available per device, potentially degrading performance. This underscores the importance of maintaining sufficiently large per-GPU batch sizes even when using DDP. To tackle this issue, we adopt the strategy implemented in OpenCLIP (Ilharco et al., 2021). Before computing the contrastive loss, each GPU (rank) gathers the feature representations from all other GPUs. These features are then re-concatenated, effectively creating a unified global batch for the loss calculation on each device. This process preserves the gradient flow for proper backpropagation while providing every GPU with a much larger and more effective set of negative samples. For clarity, we provide pseudocode for the loss computation in Alg. 1 in Sec. A.1. This implementation achieves over 5-fold faster training speed while maintaining the original model accuracy.
4 GLEAM-X: A Multi-Image MLLM Explanation Benchmark
GLEAM-X extends GLEAM-C to a multi-image reasoning setting, where the model not only predicts matches between query and reference images but also generates interpretable explanations. In practice, GLEAM-C and GLEAM-X operate as an integrated pipeline: GLEAM-C first retrieves the best-matching reference image for a given query, and GLEAM-X then verifies whether this reference image truly matches the query while providing human-interpretable explanations. In this section, we present the data composition (Sec. 4.1), annotation procedure (Sec. 4.2), and evaluation protocol (Sec. 4.3) of the GLEAM-X benchmark.
4.1 Data Composition
Our training and testing data pairs are derived from the GLEAM-C training and test sets. To maintain balance across modalities and views, we perform data sampling to ensure that the amount of training data for each modality is roughly equal, as shown in Tab. 1. In the evaluation of GLEAM-X, we also maintain this balance across modalities. The selected query images for testing include 128 for MAP, 126 for SetVL-480K, 127 for University-1652, and 123 for VIGOR. For each query image, we select one positive reference image and one negative reference image. Using the corresponding labels, we generate explanation annotations in both English and Chinese with commercial MLLMs, specifically GPT-4o (Hurst et al., 2024) and Doubao-1.5-Thinking-Vision-Pro (Volcengine, 2025).
[…]
4.2 Data Annotation
To ensure the accuracy of the test set, we manually revise the 2k test pairs. We engage 4 human experts in Remote Sensing, each holding at least a master’s degree, to perform data annotation and revision. We conduct two annotation rounds: in each round, we first verify the correctness of the explanations generated by the commercial MLLMs and correct them if necessary; if the four experts cannot determine whether a query and reference image match, we replace the pair with a suitable one from the remaining test data and perform model-assisted generation followed by manual correction. This ensures that all test pairs are valid and unambiguous, providing a reliable basis for evaluation.
4.3 Evaluation Protocol
We evaluate the MLLM outputs on the test set from two perspectives. First, we extract the predicted match labels using regular-expression parsing and compute the matching accuracy. Second, we assess the quality of the generated explanations by measuring their semantic similarity to the annotated references with Sentence-BERT (Reimers & Gurevych, 2019). This dual evaluation framework ensures a comprehensive assessment, as it captures both the correctness of structured match predictions and the semantic fidelity of explanatory outputs.
5 Experiments
In this section, we design a series of experiments to evaluate the unified GLEAM-C model (Sec. 5.1) and the GLEAM-X benchmark (Sec. 5.2). All training experiments are conducted on a cluster equipped with 10 NVIDIA RTX 4090 D GPUs, each with 48 GB of memory.
5.1 Evaluation results of GLEAM-C
For GLEAM-C, we compare the unified model with single-dataset training and evaluate the data integration strategy (Sec.5.1.1). We further assess improvements in training efficiency (Sec.5.1.2) relative to prior SOTA training methods.
Backbone Training Strategy Dataset Epochs R@1 R@5 R@10 Top-1 AP/HR
Single-dataset VIGOR 40 77.86 95.66 97.21 99.61 89.82
ConvNeXt-B-384 Sample4Geo University-1652 1 92.65 - - - 93.81
GPS + DSS SetVL-480K 40 16.86 38.95 47.71 85.40 16.86
ConvNeXt-B-384
Single-dataset
Sample4Geo
w/o GPS
VIGOR 40 76.60 95.07 96.83 99.64 88.29
University-1652 1 91.55 97.79 98.37 98.44 92.98
MAP 40 92.60 98.08 98.94 99.80 92.60
SetVL-480K 40 14.23 34.10 44.63 71.67 14.23
From-scratch
VIGOR 40 73.37 93.43 95.66 99.61 84.43
University-1652 87.08 95.22 96.59 96.79 88.96
MAP 92.52 97.81 98.79 99.80 92.52
SetVL-480K 14.33 34.30 44.77 70.75 14.33
Two-phase
(VIGOR $\rightarrow$ Merge)
VIGOR 40 75.66 94.51 96.43 99.65 86.80
University-1652 87.03 96.11 97.84 98.03 89.12
MAP 94.05 98.04 98.71 99.65 94.05
SetVL-480K 15.28 35.49 45.71 71.30 15.28
PE-Core-L14-336 Single-dataset VIGOR 40 75.53 95.48 97.18 99.67 89.46
University-1652 1 94.38 98.44 98.80 98.85 95.32
MAP 40 92.79 97.92 98.55 99.73 92.79
SetVL-480K 40 21.34 46.54 57.95 81.58 21.34
From-scratch
VIGOR 40 69.44 92.81 95.43 99.63 83.18
University-1652 93.20 97.73 98.30 98.40 94.26
MAP 93.11 98.32 98.75 99.73 93.11
SetVL-480K 22.00 46.93 58.28 81.83 22.00
Two-phase
(VIGOR $\rightarrow$ Merge)
VIGOR 40 75.96 95.46 97.17 99.68 89.44
University-1652 93.19 97.92 98.49 98.57 94.28
MAP 93.97 98.20 98.43 99.77 93.97
SetVL-480K 23.25 48.61 59.76 82.98 23.25
Table 2: Image retrieval accuracy across backbones and training strategies. For University-1652, the last evaluation metric is AP; for all other datasets, it is Hit Rate. The first three rows (highlighted in gray) are directly taken from the corresponding papers with GPS (if applicable, marked with ) and DSS sampling for reference. It can be observed that using a two-phase training strategy yields relatively stronger performance. The unified GLEAM-C model achieves results comparable to or even exceeding those obtained by single-dataset training (Sample4Geo w/o GPS-Sampling).
Dataset Mode Recall@1 Recall@5 Recall@10 Recall@top1 Hit Rate Seconds/Epoch
VIGOR DP 76.80 95.06 96.70 99.25 89.05 521
DDP 76.60 95.07 96.83 99.64 88.29 92
[…]
5.1.1 From Single Dataset to Multi-Dataset Integration
As in most CVGL studies, we first train on single-view and single-modality datasets. Training is conducted separately on the VIGOR (same), University-1652, MAP, and SetVL-480K (N = 1) datasets. Model performance is evaluated using image retrieval accuracy, with University-1652 following the Drone2Sat protocol. We adopt ConvNeXt-B-384 (88.6M) and PE-Core-L14-336 (0.32B) as the image encoders. The results for single-dataset training are presented in Tab. 2. Since some datasets lack GPS metadata, we do not use GPS sampling in all our model training, which may result in a slight decrease in accuracy. For reference, we also report the original metrics from the respective papers, with GPS (if applicable) and DSS sampling (Deuser et al., 2023b; Wu et al., 2024a).
To train the unified model, we compare two data integration strategies in our work. The first strategy trains on mixed data directly from the backbone’s initial weights, while the second strategy starts training from a VIGOR checkpoint obtained through single-dataset training. We use the data sampling ratios shown in Tab. 1, and both strategies are trained for 40 epochs. We provide the results in Tab. 2 and share the following observations:
1) Directly training on the mixed data degrades the performance on VIGOR. However, continuing training the unified model from a VIGOR checkpoint yields better results and improves accuracy on MAP and SetVL-480K. This highlights the importance of endowing the model with basic CVGL capability before performing mixed-data training, and also suggests that learning across different views and modalities may mutually reinforce each other.
2) Compared with single-dataset training, our two-phase training strategy yields the foundational GLEAM-C model, which attains comparable or superior performance across all evaluated datasets.
3) The larger PE-Core-L14-336 model achieves overall better performance compared to ConvNeXt-B-384 after training. Additionally, although University-1652 requires only a single epoch of training, we train for 40 epochs due to the demands of mixed-data training; under this setting, smaller models may suffer from overtraining or instability, whereas larger models remain largely unaffected.
5.1.2 Training Efficiency Evaluation
We evaluate training efficiency on the VIGOR dataset by comparing DP and DDP strategies. Following the experiment settings of Sample4Geo (Deuser et al., 2023b), we adopt the ConvNeXt-B-384 backbone and use the original Sample4Geo code for DP training (without GPS-based sampling). As shown in Tab. 3, both approaches achieve comparable image retrieval accuracy across all Recall@k and Hit Rate metrics. However, DDP demonstrates a substantial (5-fold) speed advantage, reducing the time per training epoch from 521 seconds to 92 seconds, highlighting its effectiveness for accelerating large-scale CVGL model training.
5.2 Evaluation results on GLEAM-X
Model Language MAP SetVL-480K University-1652 VIGOR Avg Acc
GPT-4o 85.55 63.89 93.70 81.71 81.25
Doubao-1.5 91.80 74.21 60.24 83.74 77.48
Qwen2.5-VL-3B-Instruct 55.08 57.94 51.57 65.04 57.34
   +label only 96.88 84.13 97.64 92.28 92.76
   +GPT-4o 94.53 78.17 99.61 89.02 90.38
   +Doubao-1.5 EN 95.31 79.37 96.46 87.80 89.78
Model Language MAP SetVL-480K University-1652 VIGOR Avg Acc
GPT-4o 85.55 70.63 96.06 82.11 83.63
Doubao-1.5 89.45 78.97 66.93 81.30 79.17
Qwen2.5-VL-3B-Instruct 51.17 51.19 50.00 49.59 50.50
   +label only 96.88 84.13 97.64 92.28 92.76
   +GPT-4o 95.31 76.59 99.61 89.02 90.18
   +Doubao-1.5 ZH 94.53 75.00 94.88 86.59 87.80
Table 4: Matching accuracy comparison of different models on GLEAM-X in English and Chinese. Results are reported on MAP, SetVL-480K, University-1652, and VIGOR datasets, along with the overall average accuracy. Label-only supervision and LLM-explanation supervision (+GPT-4o, +Doubao-1.5) significantly improve the performance of Qwen2.5-VL-3B-Instruct.
Model Language 0.0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1.0 Avg Sim
GPT-4o 0.0% 0.1% 0.8% 21.6% 77.5% 0.8349
Doubao-1.5 0.0% 0.0% 0.2% 33.5% 66.3% 0.8171
Qwen2.5-VL-3B-Instruct 0.8% 0.0% 2.0% 59.3% 37.9% 0.7645
   +label only 99.9% 0.1% 0.0% 0.0% 0.0% 0.0981
   +GPT-4o 0.0% 0.0% 0.1% 22.4% 77.5% 0.8405
   +Doubao-1.5 EN 0.0% 0.0% 0.5% 36.0% 63.5% 0.8122
Model Language 0.0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1.0 Avg Sim
GPT-4o 1.6% 5.0% 2.3% 50.5% 40.7% 0.7403
Doubao-1.5 0.0% 0.0% 4.5% 66.7% 28.9% 0.7495
Qwen2.5-VL-3B-Instruct 4.2% 0.1% 2.9% 58.5% 34.3% 0.7399
   +label only 92.1% 7.9% 0.0% 0.0% 0.0% 0.1476
   +GPT-4o 0.0% 0.0% 0.3% 32.5% 67.2% 0.8239
   +Doubao-1.5 ZH 0.4% 0.0% 1.1% 46.6% 51.9% 0.7924
Table 5: Similarity score distribution (by Sentence-BERT) for different models on GLEAM-X in English and Chinese. Columns show percentages in each similarity range (0.0–0.2, …, 0.8–1.0) and the last column reports Avg Sim. LLM-explanation supervision (+GPT-4o, +Doubao-1.5) improves alignment with ground-truth, while label-only supervision fails to generate explanations.
We evaluate both commercial models and open-source models on the GLEAM-X benchmark. For commercial models, we regenerate answers on the test set using GPT-4o and Doubao-1.5-Thinking-Vision-Pro without providing ground-truth labels. For open-source models, we compare three variants of Qwen2.5-VL-3B-Instruct: the original model, the version fine-tuned with LLM-explanation supervision, and a version trained solely with label-only supervision. Specifically, for the explanation-based variant, we follow the procedure in Sec. 4.1, concatenating labels and explanations during training. We train separately on the 200k explanations generated by GPT-4o and Doubao-1.5 for 1 epoch. Conversely, the label-only variant uses only labels (e.g., [[1]]) for training.
After inference on the test set, we decouple matching predictions and explanations through regularization to isolate their respective contributions. To provide a comprehensive evaluation of GLEAM-X, we perform detailed analyses from three perspectives: matching accuracy (Sec.5.2.1), semantic accuracy (Sec.5.2.2), and positive/negative sample analysis (Sec. 5.2.3).
5.2.1 Matching Accuracy
The matching accuracy results are presented in Tab. 4. Commercial models achieve a relatively strong performance of around 80%, with GPT-4o outperforming Doubao-1.5 on average. In contrast, the original Qwen2.5-VL-3B-Instruct model exhibits very low accuracy, particularly in the Chinese setting, where its accuracy is only 50.50%. After training, all model variants obtain accuracy surpassing that of the commercial MLLMs. Notably, the label-only supervision model achieves the highest accuracy, likely because its training targets are extremely short (e.g., [[0]] or [[1]]), which reduces the learning difficulty for the model.
5.2.2 Semantic Analysis of Explanations
Tab. 5 reveals the similarity distributional differences across models. Commercial MLLMs concentrate heavily in the high-similarity range (0.8–1.0) for English, but shift toward 0.6–0.8 in Chinese. The original Qwen2.5-VL-3B-Instruct places a substantial proportion of predictions in the mid-similarity bins (0.6–0.8) for both languages, indicating weaker semantic alignment with ground-truth explanations. Fine-tuning with LLM-explanation supervision shifts the distribution upward, with both +GPT-4o and +Doubao-1.5 variants peaking in the 0.8–1.0 range. By contrast, the label-only supervision model collapses almost entirely into the lowest bin (0.0–0.2), confirming its inability to generate semantically meaningful explanations despite improvements in matching accuracy.
5.2.3 Analysis of Positive and Negative Pair Samples
Model Language Pos Acc Neg Acc Pos-Neg Pos Sim Neg Sim Pos-Neg
GPT-4o 70.83 91.67 -20.84 0.8296 0.8401 -0.0105
Doubao-1.5 66.87 88.10 -21.23 0.8100 0.8241 -0.0141
Qwen2.5-VL-3B-Instruct 23.81 90.87 -67.06 0.7616 0.7675 -0.0059
   +label only 93.85 91.67 2.18 0.0851 0.1111 -0.0260
   +GPT-4o 93.25 87.50 5.75 0.8399 0.8411 -0.0012
   +Doubao-1.5 EN 90.08 89.48 0.60 0.8048 0.8195 -0.0147
Model Language Pos Acc Neg Acc Pos-Neg Pos Sim Neg Sim Pos-Neg
GPT-4o 73.02 94.25 -21.23 0.7337 0.7468 -0.0131
Doubao-1.5 76.19 82.14 -5.95 0.7435 0.7555 -0.0120
Qwen2.5-VL-3B-Instruct 1.19 99.80 -98.61 0.7265 0.7533 -0.0268
   +label only 93.85 91.67 2.18 0.1471 0.1481 -0.0010
   +GPT-4o 90.87 89.48 1.39 0.8211 0.8267 -0.0056
   +Doubao-1.5 ZH 84.52 91.07 -6.55 0.7858 0.7989 -0.0131
Table 6: Positive and negative sample accuracy (Pos/Neg Acc) and similarity (Pos/Neg Sim) for different models on GLEAM-X in English and Chinese. The Pos-Neg columns report the difference between positive and negative samples. Vanilla Qwen2.5-VL-3B-Instruct shows a strong bias toward predicting non-matching pairs, while LLM-explanation supervision (+GPT-4o, +Doubao-1.5) mitigates this bias and improves alignment with ground-truth explanations.
To further analyze model behavior on positive and negative samples, we examine both matching accuracy and semantic similarity in Tab. 6. Vanilla MLLMs, including both commercial models and Qwen2.5-VL-3B-Instruct, exhibit a bias toward negative samples. In particular, Qwen2.5-VL-3B-Instruct achieves very low positive accuracy while maintaining high negative accuracy, indicating that the original model tends to predict image pairs as non-matching. Fine-tuning with LLM-explanation supervision (+GPT-4o, +Doubao-1.5) substantially mitigates this bias, raising positive accuracy to over 90% in English and over 84% in Chinese, while still keeping negative accuracy high. Correspondingly, positive and negative similarity are more balanced. Overall, LLM-explanation supervision effectively improves both prediction balance and explanation quality across languages. In contrast, the label-only supervision variant achieves high positive accuracy but fails to generate semantically meaningful explanations, as evidenced by extremely low similarity scores.
6 Conclusion
In this paper, we systematically address the challenges of cross-view geo-localization by making several key contributions. First, we provide a detailed analysis of the CVGL problem, highlighting the benefits of a unified framework for integrating multiple views and modalities, and emphasizing the importance of interpretability and system transparency in practical applications. Second, we propose GLEAM-C, a foundational model that aligns diverse viewpoints and modalities with satellite imagery, improving training efficiency and achieving high accuracy through a two-phase training strategy. Third, we introduce GLEAM-X, a bilingual benchmark that enables explainable reasoning, allowing the generation of human-interpretable explanations for image correspondences and enhancing model transparency. GLEAM-C and GLEAM-X can be integrated into a unified pipeline that combines accurate correspondence prediction with interpretable explanations, improving both robustness and trustworthiness of CVGL systems. We hope that our work will contribute to advancing research and development in this field of study.
Appendix A Appendix
A.1 Pseudocode for the DDP Loss Computation
We provide the pseudocode for DDP loss computation below.
Input: Image features $\mathbf{f}_{1},\mathbf{f}_{2}\in\mathbb{R}^{B\times d}$, logit scale $\alpha$, world size $W$, rank $r$
Output: Contrastive loss $\mathcal{L}$ on rank $r$
1ex
1Normalize features: $\mathbf{f}_{1}\leftarrow\mathrm{normalize}(\mathbf{f}_{1})$, $\mathbf{f}_{2}\leftarrow\mathrm{normalize}(\mathbf{f}_{2})$;
2Gather features from all processes:
$\{\mathbf{f}_{1}^{(w)}\}_{w=1}^{W}\leftarrow\mathrm{AllGather}(\mathbf{f}_{1}),\quad\{\mathbf{f}_{2}^{(w)}\}_{w=1}^{W}\leftarrow\mathrm{AllGather}(\mathbf{f}_{2})$
3Concatenate features, excluding local rank duplicates:
$\mathbf{F}_{1}\leftarrow[\mathbf{f}_{1}]\cup\{\mathbf{f}_{1}^{(w)}\mid w\neq r\},\quad\mathbf{F}_{2}\leftarrow[\mathbf{f}_{2}]\cup\{\mathbf{f}_{2}^{(w)}\mid w\neq r\}$
4Compute logits:
$\mathbf{Z}_{1}=\alpha\cdot\mathbf{F}_{1}\mathbf{F}_{2}^{\top},\quad\mathbf{Z}_{2}=\mathbf{Z}_{1}^{\top}$
5Construct global labels:
$\mathbf{y}=[0,1,\dots,N-1],\quad N=\text{rows}(\mathbf{Z}_{1})$
6Compute loss:
$\mathcal{L}=\tfrac{1}{2}\big(\mathrm{CE}(\mathbf{Z}_{1},\mathbf{y})+\mathrm{CE}(\mathbf{Z}_{2},\mathbf{y})\big)$
return $\mathcal{L}$;
Algorithm 1 Distributed InfoNCE Loss (DDP)
A.2 Details of Human Annotators
In the Data Annotation stage, we hire 4 human experts in Remote Sensing, each holding at least a master’s degree. We pay them a labeling fee of \$20 per hour.
A.3 Hyper-parameters
A.3.1 GLEAM-C
Here we provide the training hyper-parameters of GLEAM-C using PE-Core-L14-336 ViT on the second training phase.
Parameter Value
Model PE-Core-L14-336
Epochs 40
Batch Size (Total) 300
Learning Rate 1e-4
Scheduler Cosine
Warmup Epochs 1
Label Smoothing 0.1
Neighbour Select 64
Neighbour Range 128
Prob Rotate 0.75
Prob Flip 0.5
Table 7: Training hyper-parameters of GLEAM-C using PE-Core-L14-336.
A.3.2 GLEAM-X
Here we provide the hyper-parameters of training Qwen2.5-VL-3B-Instruct on GLEAM-X.
Parameter Value
Learning Rate 5e-5
Batch Size (Total) 60
Epochs 1
Weight Decay 0.1
Warmup Ratio 0.05
Max Length 2048
Min/Max Pixels 256/1296
Gradient Checkpointing True
Precision BF16
Table 8: Hyper-parameters of training Qwen2.5-VL-3B-Instruct on GLEAM-X.
A.4 More Data Examples of GLEAM-X.
Figure 3: Street map and ground photograph samples.
Figure 4: Panoramic view and UAV imagery samples.
A.5 More Qualitative Results on GLEAM-X Benchmark
Figure 5: Sample on the VIGOR test set (Chinese scenario). The gray English text is a direct translation of the Chinese response.
Figure 6: Sample on the University-1652 test set (English scenario).